### v1

# Breaking the SFT Plateau: Multimodal Structured Reinforcement Learning for Chart-to-Code Generation

Abstract
While reinforcement learning (RL) has proven highly effective for general reasoning in vision-language models, its application to tasks requiring in-depth understanding of information-rich images and the generation of structured outputs remains underexplored. Chart-to-code generation is a prime example of this challenge, as it demands complex reasoning over visual charts to generate structured code. Therefore, supervised fine-tuning (SFT) alone is often insufficient, highlighting a critical need for effective RL strategies that appropriately reward the structured outputs. In this paper, we systematically investigate the performance plateau in SFT through large-scale experiments and propose Multimodal Structured Reinforcement Learning (MSRL) for chart-to-code generation, which substantially breaks through this plateau. To this end, we first construct the largest training corpus to date, containing 3 million chart-code pairs. To mitigate the simplistic patterns of prior synthetic data, we curate real-world tables from arXiv papers as the data source. Despite reaching a state-of-the-art performance, our experiments demonstrate that simply scaling the SFT data eventually hits a performance plateau where further increases yield negligible improvements. To break this plateau, our MSRL method leverages a multi-granularity structured reward system that uses multimodal textual and visual feedback to guide the RL training. At the textual level, rule-based rewards validate the fidelity of fine-grained code details. At the visual level, a model-based reward is derived by first rendering the generated code into an image and then employing an evaluator model to assess its structural similarity to the ground-truth chart. To ensure training stability, we implement this multimodal reward within a two-stage curriculum, where the model initially trains on textual rewards alone before progressing to the combined textual and visual signals. Experimental results demonstrate that our MSRL framework significantly breaks the SFT plateau, improving the high-level metrics by 6.2% and 9.9% on ChartMimic and ReachQA benchmarks respectively. Notably, our method significantly outperforms all methods in the chart domain and achieves competitive performance with advanced closed-source models for the first time.
1 Introduction
Large language models (LLMs) have demonstrated impressive reasoning capabilities on complex textual problems, including code generation and mathematical problem-solving OpenAI (2025b); Guo et al. (2025). Pivotal to this success is the implementation of reinforcement learning (RL) to optimize the capacities of LLM to generate step-by-step reasoning Wei et al. (2022) and verify final answers Shao et al. (2024). While multimodal large language models (MLLMs) have achieved impressive performance on established visual reasoning tasks like visual question answering Zhang et al. (2024a); Huang et al. (2025), their core reasoning capabilities remain limited, especially when tasks requiring processing information-dense images like charts Xu et al. (2024); Chen et al. (2025a), and generating structured outputs like code Xia et al. (2024); Zhao et al. (2025).
Chart-to-code generation, a complex task characterized by dense visual inputs and structured code outputs, has recently garnered significant attention Shi et al. (2024); Zhang et al. (2024c). Although previous works have attempted to solve this task by using methods like Supervised Fine-Tuning (SFT) Zhang et al. (2024b); Zhao et al. (2025) or Direct Preference Optimization (DPO) Rafailov et al. (2023); Zhang et al. (2025) on synthetic data, their performance remains limited due to the simplistic patterns of the corpus and ineffective generalization strategies. Therefore, a promising direction is to leverage more realistic data and robust strategies to enhance model performance.
Figure 1: The SFT plateau and RL performance gain from our experiments. The left figure illustrates that scaling SFT data from 200k to 2.8M leads to a performance plateau after exceeding 2M data points. The right figure denotes the performance gain from our proposed MSRL training strategy.
In this work, we first systematically investigate the SFT performance plateau through experiments at various data scales. Based on this analysis, we propose a multimodal structured reinforcement learning (MSRL) strategy designed to optimize the model at multiple granularities. Specifically, to investigate the performance plateau of SFT, we first construct the largest training corpus to date, which covers 24 distinct chart types and comprises 3 million chart-code pairs. To ensure the charts reflect realistic distributions and avoid monotonous trends, we design a two-stage generation pipeline that curates high-quality tables from arXiv papers and subsequently leverages these data to generate diverse and complex plot codes utilizing LLMs. Unlike previous works that apply RL directly after SFT Ni et al. (2025); Chen et al. (2025b), our approach first investigates the performance plateau of SFT models. Establishing this baseline is crucial for isolating the true effectiveness of the MSRL strategy from gains that merely compensate for an unsaturated SFT model. To this end, we perform SFT on models with various data scales and reveal that simply increasing the SFT data volume leads to a performance plateau, which cannot be overcome by data quantity alone. To break through this plateau, we propose the MSRL training strategy, which incorporates multi-granularity reward functions from both textual and visual aspects to optimise the model. For the textual aspect, to handle format diversity in code generation, we first filter our training corpus to enforce a consistent structure, making it suitable for information extraction. Building on this standardized data, we develop a comprehensive reward function that verifies code correctness from five key perspectives. However, a reward function based solely on text is often insufficient, as it struggles to capture the overall structures of charts. To address this limitation, we introduce a visual reward mechanism. The core of this method is a render-and-compare process that involves rendering the generated code into an image and then employing an MLLM to score the visual similarity between the rendered image and the original input chart. By leveraging its multi-granularity reward function, MSRL optimizes both global context and fine-grained details. We evaluate models on various chart-to-code benchmarks Shi et al. (2024); He et al. (2024). The results demonstrate that it establishes a new state-of-the-art, as it outperforms all previous open-source models and rivals the performance of powerful proprietary MLLMs. In summary, the main contributions of this work are as follows:
  • We propose an MSRL training strategy that employs a multi-level reward function, which combines a rule-based component for assessing textual code correctness with a model-based evaluator for evaluating the visual fidelity of the rendered chart. Benefits from the MSRL training strategy, our proposed model achieves state-of-the-art (SOTA) performance compared to all other open-source MLLMs.
  • We construct the first large-scale chart-to-code training corpus, generated using real-world tables as a data source. After applying several filtering methods, we curate a final dataset of 2.8 million samples for SFT and 33 thousand for RL.
  • Through SFT experiments at various data scales, we identify the performance bottleneck, confirming that merely increasing SFT data quantity is insufficient to improve performance.
2 Related Work
2.1 Chart MLLMs
Chart understanding is a crucial research area encompassing both low-level tasks, such as data extraction Liu et al. (2022), and high-level tasks like question answering Masry et al. (2022) and summarization Kantharaj et al. (2022). Recent work has focused on training MLLMs on extensive, chart-specific datasets to enhance their understanding capabilities Xia et al. (2024); Zhang et al. (2024b), leading to superior performance across various chart-related tasks. For example, ChartMoE Xu et al. (2024) proposes a mixture-of-experts (MoE) structure for multi-task pretraining. More recently, leveraging reinforcement learning to enhance reasoning capacities has garnered significant interest. Both Chart-R1 Chen et al. (2025a) and BigCharts-R1 Masry et al. (2025) optimize MLLMs using Chain-of-Thought (CoT) Wei et al. (2022) reasoning data and reinforcement learning with verified rewards (RLVR) Shao et al. (2024).
2.2 Reasoning MLLMs
The success of large reasoning models (LRMs) like DeepSeek-R1 Guo et al. (2025) spurs significant research into enhancing LLM reasoning through reinforcement learning with rule-based rewards Shao et al. (2024). This paradigm extends to the vision-language domain, where numerous works aim to improve the long-chain reasoning capabilities of MLLMs Shen et al. (2025); Wang et al. (2025b). Following this trend, initial efforts such as Vision-R1 Huang et al. (2025) and R1-OneVision Yang et al. (2025) adapt the Group Relative Policy Optimization (GRPO) algorithm with multimodal reasoning data to enable long-form reasoning in VLMs. Subsequent works, including MMEureka Meng et al. (2025b) and R1-Zero Liu et al. (2025), further advance this area by introducing improved reinforcement learning strategies for visual long-term reasoning. Recently, approaches like Point-RFT Ni et al. (2025) utilize task-specific thinking frameworks to enhance reasoning capabilities further.
2.3 MLLMs For Code
In the rapidly advancing research area of multimodal code generation, foundational work has focused on establishing benchmarks to assess the capabilities of MLLMs. These benchmarks evaluate performance on tasks such as generating execution code for solving visual problems Li et al. (2024); Wang et al. (2025a) and HTML code from webpage screenshots Si et al. (2024); Yun et al. (2024); Xiao et al. (2024). Among these emerging tasks, chart-to-code generation has attracted significant interest due to the complexity of its visual inputs. This task challenges an MLLM to generate plotting code that accurately reproduces the chart image. Recently, several benchmarks have been proposed to evaluate MLLMs in this context, assessing a range of capabilities. For instance, benchmarks such as ChartMimic Shi et al. (2024), Plot2Code Wu et al. (2024), and ChartX Xia et al. (2024) assess chart-to-code generation capabilities.
3 Methods
Figure 2: The data generation pipeline and our proposed MSRL framework. (a) Our pipeline prompts Gemini-2.5-Flash with tables from arXiv papers and example codes to generate plotting code. After execution and filtering, the chart-to-code dataset of 3M pairs is obtained. We then sample 130k of these pairs and apply three-stage filters (chart, data, vision) to curate a final, high-quality dataset of 33k examples for RL. (b) The framework of our proposed MSRL strategy. The textual reward is derived from a rule-based evaluation of the generated code across five distinct dimensions. An MLLM quantifies the visual reward based on the rendered image.
3.1 Training Corpus Construction
While several chart-to-code generation datasets have been proposed, they are generally constrained by a reliance on purely synthetic data and limited scale, resulting in charts that exhibit simplistic trends and lack diversity. To overcome these challenges, we construct our dataset by crawling real-world tables from arXiv papers and then leveraging LLMs to generate the corresponding plotting code. We ensure data diversity and quality through several strategies, such as using high-quality seed code and filtering out non-executable code.
Through this process, we construct the largest training corpus to date, comprising 3 million chart-code pairs, to investigate the performance limits of SFT. However, previous research Chen et al. (2025a); Huang et al. (2025) shows that separating the SFT and RL training corpora prevents overfitting to the SFT data format and enhances the exploration capacity of RL. To curate a high-quality dataset for RL, we begin with a candidate corpus of 130 thousand samples from the overall training corpus. These candidate data samples are then subjected to a two-stage filtering strategy. The first filtering stage focuses on the code content, filtering based on chart and data type to ensure code diversity and that the output code is suitable for data extraction. Specifically, the chart types are filtered according to plot functions. To handle complex charts containing multiple chart types, we employ a tree-structured parsing method to progressively identify each type present. The data values are filtered according to the data definition format, retaining only data structured as one-dimensional arrays and non-nested dictionaries to facilitate extraction during the RL stage. This step narrows the candidate corpus to 45 thousand samples. The second stage addresses visual quality, employing GPT-4.1-mini OpenAI (2025a) as a visual LLM judge to filter for high-fidelity chart images. This process yields the final RL training corpus of 33 thousand high-quality samples. The remaining samples are utilized for SFT.
3.2 Supervised Fine-tuning
Firstly, we perform SFT on Qwen2.5-VL as the base model using our generated chart-to-code dataset. To explore the effectiveness of data scaling and identify the SFT plateau, we fine-tune our model on six distinct training corpus sizes, created by partitioning our main dataset into subsets of 200k, 400k, 800k, 1.6M, 2M and 2.8M samples. The model is trained by minimizing the standard autoregressive objective, which is the negative log-likelihood of the target sequence. As demonstrated in Section 4.3 and Figure 1, merely scaling the SFT data volume ultimately results in a performance plateau. Increasing the SFT dataset from 2M to 2.8M samples results in no further performance gains.
3.3 Multimodal Structured Reinforcement Learning
Although SFT can significantly enhance the chart-to-code capabilities of MLLMs, our experiments reveal that this approach ultimately reaches a performance plateau, likely due to the limitation that SFT restricts the model’s capacity to restore detailed information. This limitation stems from the core drawback of the SFT objective that it treats every token in the target sequence with uniform importance. Therefore, SFT is ineffective in optimizing plotting code generation, as the plotting code comprises largely of boilerplate snippets such as plt.plot, with critical information like specific data values or styling parameters appearing infrequently. Consequently, employing RL with a reward function designed to prioritize the accuracy of such critical content presents a promising approach for enhancing code generation capacity. To this end, we propose a two-stage training strategy to stabilize the RL process, utilizing multi-granularity reward functions that combine both textual and visual feedback.
Textual Reward While the RLVR method provides a powerful mechanism to guide models by measuring the correctness of detailed information, its practical application is challenged by the stylistic diversity of generated code, particularly in data definitions and function calls. This stylistic heterogeneity leads to difficulty in comprehensively extracting and parsing the critical information for reward computation. To mitigate this, we introduce a code normalization step that maps each generated output to a canonical representation before calculating the reward. This process ensures the textual reward function is invariant to syntactic variations, providing a more suitable method for RLVR. Specifically, the textual reward is multifaceted, consisting of a binary score for execution success and a more granular rule-based accuracy score. This accuracy metric is a weighted average derived from evaluating key aspects of the generated code. We assess data values using soft value matching, chart types with hard string matching, layout with hard value matching and elements like titles and labels via edit distance. The soft value matching allows a relative error tolerance of $\pm$5%. The final textual reward is the addition of execution and a rule-based accuracy score.
Visual Reward However, our analysis reveals that purely textual rewards generally focus on fine-grained details, while the overall structure of the rendered chart image is not considered. The chart-to-code task requires generating code that replicates the overall style of charts. To this end, we propose visual reward feedback. Specifically, the generated code is executed to render a chart image, and we then utilize MLLMs to score the visual similarity between the generated chart and the input. This score is then normalized to serve as the visual reward. For the visual reward, we employ Qwen2.5-VL-72B Bai et al. (2025) as the evaluation model, which compares the chart image rendered by the generated code against the ground-truth chart across six key aspects: chart type, layout, text content, data, style, and clarity. The resulting scores are then normalized to form the final reward. Crucially, code that fails to render an image receives a reward of 0. The framework of our proposed MSRL is denoted in Figure 2.
Two-stage RL To ensure training stability, our two-stage RL strategy first optimizes the model using only a textual reward. In the second stage, we introduce a hybrid reward, which combines the textual reward with a visual one, to fine-tune the model for visual fidelity. To properly balance the different training objectives, the total reward $R$ for a completed trajectory is calculated as a weighted sum of the following three components:
$R=w_{t}R_{\text{text}}+w_{v}R_{\text{vis}}+w_{e}R_{\text{exec}}$ (1)
where $w_{t}$, $w_{v}$, and $w_{e}$ are hyperparameters that balance the contribution of each component.
GRPO We adapt the Group Relative Policy Optimization (GRPO) Shao et al. (2024) algorithm for RL. For each input $(x,y)$, the policy $\pi_{\theta}$ samples a group of $G$ candidate responses $\{o_{i}\}_{i=1}^{G}$.
$\mathcal{J}_{GRPO}(\theta)=\mathbb{E}_{\begin{subarray}{c}(x,y)\sim\mathcal{D}_{\text{CoT}},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid x)\end{subarray}}$ $\biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\biggl{(}\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)}A_{i},$ (2)
$\operatorname{clip}\left(\frac{\pi_{\theta}(o_{i}\mid x)}{\pi_{\theta_{\text{old}}}(o_{i}\mid x)},1-\varepsilon,1+\varepsilon\right)A_{i}\biggr{)}]$
where $\varepsilon$ is a hyperparameter, and $\pi_{\theta}$, $\pi_{\theta_{\text{old}}}$ are the optimized model and the old policy model. The group-normalized advantage for the $i$-th response is:
$A_{i}=\frac{r_{i}-\operatorname{mean}\left(\left\{r_{1},r_{2},\cdots,r_{G}\right\}\right)}{\operatorname{std}\left(\left\{r_{1},r_{2},\cdots,r_{G}\right\}\right)}$ (3)
4 Experiment
4.1 Baselines and Benchmarks
We evaluate MSRL against three categories of state-of-the-art models. The first category comprises general-domain, open-source Multimodal Large Language Models (MLLMs): InternVL2 (8B, 26B) Chen et al. (2024), Qwen2-VL (7B, 72B) Wang et al. (2024), and Qwen2.5-VL-7B Bai et al. (2025). The second includes proprietary models: GeminiProVision Team et al. (2023), Claude-3-opus Anthropic (2024), GPT-4V OpenAI (2023), and GPT-4o OpenAI (2024). The third category consists of chart-specific MLLMs: ChartLlama Han et al. (2023), Tinychart Zhang et al. (2024b), ChartVLM Xia et al. (2024), Chart2Code Zhang et al. (2025) and ChartCoder Zhao et al. (2025).
All models were evaluated in a zero-shot setting on two distinct benchmarks: ChartMimic Shi et al. (2024) and ReachQA He et al. (2024). For ReachQA, we adopt the evaluation setting of Chart2Code Zhang et al. (2025), using the 500 plotting scripts from its test set. For ChartMimic, we utilize the 600 examples from the latest version of the Direct Mimic task.
Model Params ChartMimic ReachQA
Exec.Rate Low-Level High-Level Exec.Rate Low-Level High-Level
Proprietary
GeminiProVision - 68.2 53.8 53.3 74.0 67.0 67.8
Claude-3-opus - 83.3 60.5 60.1 89.0 51.7 61.1
GPT-4V - 91.2 76.4 78.9 88.0 69.5 78.6
GPT-4o - 93.2 79.0 83.5 92.8 81.8 84.0
Open-Source General-Domain
Qwen2-VL-7B 7B 67.0 32.9 35.0 55.4 22.6 29.3
Qwen2.5-VL-7B 7B 73.2 44.6 41.6 62.2 36.9 37.6
InternVL2-8B 8B 61.8 34.4 38.9 50.8 24.1 24.2
InternVL2-26B 26B 69.3 41.4 47.4 55.4 29.0 28.8
Qwen2-VL-72B 72B 73.3 54.4 50.9 77.2 50.0 48.1
Open-Source Chart-Domain
ChartLlama 13B 57.5 24.8 28.1 54.8 11.1 8.1
TinyChart 3B 42.5 26.3 25.9 34.4 11.6 11.2
ChartVLM-L 14B 19.5 15.8 13.9 8.2 2.1 3.9
Chart2Code 7B 62.1 42.9 33.3 63.6 52.3 49.7
ChartCoder 7B 91.4 72.5 74.0 83.8 67.9 69.4
MSRL-SFT 7B 93.2 73.0 77.6 92.2 78.6 80.0
MSRL 7B 96.5 78.6 83.8 98.2 86.1 89.9
Table 1: Evaluation results of various baseline models. $\ddagger$} denotes the updated result with code_pass constraints reported in its repository.
4.2 Implementation Details
We utilize 2.8M chart2code data for SFT and 33k curated high-quality chart2code data for RL. MSRL-SFT employs Qwen2.5VL-7B-Instruct as the baseline model and LLaMA-Factory Zheng et al. (2024) for SFT. The SFT stage is trained for one epoch with a learning rate of 1e-5 and a batch size of 32. The training process takes 60 hours on 16 H800 GPUs. Built upon the SFT model, MSRL undergoes a two-stage RL process using the MM-EUREKA framework Meng et al. (2025a). MSRL consists of two RL stages. Stage one utilizes 22k data pairs and a textual-only reward ($w_{t}=1$, $w_{v}=0$, $w_{e}=0.5$). Stage two uses the remaining 11k pairs with a hybrid reward ($w_{t}=0.5$, $w_{v}=0.5$, $w_{e}=0.5$). The textual reward components are weighted as $w_{d}=0.4$, $w_{c}=0.3$, $w_{l}=0.1$, $w_{t}=0.1$, $w_{lbl}=0.1$. We deploy Qwen2.5-VL-72B-Instruct as the evaluation model with max_model_len set to 12,288 for calculating visual rewards. For both RL stages, we train for one episode with a learning rate of 1e-6 and a batch size of 128. We set the generation temperature to 1, generate 8 rollouts per sample, and omit the KL divergence from the loss computation. The first and second stages required 10 and 24 hours of training, respectively, on 24 H800 GPUs.
Model ChartMimic ReachQA
Text Layout Type Color Text Layout Type Color
GPT-4o 81.5 89.8 77.3 67.2 84.4 91.2 81.3 70.5
Qwen2-VL-72B 43.2 80.5 54.6 39.4 41.0 52.2 59.1 47.8
Qwen2.5-VL-7B 37.7 65.6 42.6 32.4 29.6 40.9 44.2 32.8
ChartCoder 65.9 85.4 72.3 66.4 58.2 80.0 70.4 63.2
MSRL 78.0 93.0 75.3 68.2 87.5 96.6 83.1 77.2
Table 2: Detailed results of low-level scores on the ChartMimic Direct Mimic and ReachQA benchmarks. The ChartCoder performance metrics have been corrected due to a flawed evaluation setup.
4.3 Main Results
Our proposed MSRL establishes a new state-of-the-art among all open-source MLLMs, as detailed in Table 1. On the ChartMimic benchmark Shi et al. (2024), it achieves an execution rate of 96.5, a low-level score of 78.6, and a high-level score of 83.8. Notably, its performance is not only superior to its open-source models but also comparable to that of GPT-4o, a significantly larger proprietary model. On the ChartMimic benchmark, MSRL outperforms ChartCoder, the previous leading Chart-domain MLLM specifically designed for chart-to-code generation. Unlike other works that yield suboptimal performances from fine-tuning open-source models, our initial MSRL-SFT model already establishes a new state-of-the-art among open-source models. However, MSRL-SFT still falls short of GPT-4o on both low-level and high-level performance metrics. Building upon this strong foundation, the MSRL strategy further boosts the performance, especially in high-level scores, demonstrating the effectiveness of our integrated approach. The performance gap between SFT and RL models is emphasized in Figure 1.
We also conduct a detailed evaluation of low-level performance metrics, with the results presented in Table 2. Notably, MSRL outperforms all open-source models across low-level evaluation metrics and surpasses GPT-4o on the majority of evaluation metrics, which shows the strong capacity of MSRL in chart-to-code generation. As the GPT-4o is much larger than MSRL, the results demonstrate the effectiveness of our proposed methods.
SFT Plateau To investigate and demonstrate the effectiveness and efficiency of RL, we utilize various data scales for SFT to explore the upper limit of it. As illustrated in Figure 1, the results reveal a clear trend, whereby model performance improves sharply as the data size increases to 400k samples, followed by a period of slower growth up to the 1.6M samples. Beyond this data scale, performance saturates and reaches a distinct plateau, showing negligible growth with any further increase in data scale. Based on this trend, even if the data is multiplied, the benefits brought by SFT will be minimal. Applying our proposed MSRL method yields significant performance gains across all metrics. On the ChartMimic benchmark, we observe an average improvement of 5%, with notable increases in execution rate as well as in both low-level and high-level scores.
4.4 Ablation Study
In this subsection, we perform various ablation studies to explore and verify the proposed two-stage training stages and multi-granularity reward functions.
SFT RL ChartMimic
Exec.Rate Low-Level High-Level
73.16 44.58 41.55
93.17 73.02 77.61
93.83 65.60 62.29
97.00 78.06 82.69
Table 3: Ablation studies on SFT and RL stages.
SFT/RL Ablation To understand the impact of different training settings, we conduct ablation studies on both the SFT and RL training stages. In this ablation study, we use only textual rewards to ensure the results are not influenced by our improved RL algorithm, allowing us to demonstrate the interplay between SFT and RL. Using Qwen2.5-VL-7B-Instruct as the baseline model, the results are presented in Table 3. As the RL-only model underperforms on low-level and high-level scores compared to the SFT-only approach, only the execution rate is slightly higher due to the inclusion of an execution-based reward. Furthermore, the significantly larger performance gain from RL on a suboptimal baseline, compared to an SFT model, indicates that this improvement is partly attributable to the unsaturated state of baseline models. The performance gain on a saturated SFT model demonstrates that our proposed multi-granularity reward function can surpass the upper limit of SFT, achieving significantly superior performance. Figure 3 compares the reward and execution rate for the baseline and SFT models during the RL training, showing that applying RL to the SFT model achieves a much higher textual reward.
Figure 3: Comparison of textual reward and execution rate changes between baseline and SFT models during the RL stage.
Figure 4: Comparison of reward gains during RL training with various reward settings.
Rewards ChartMimic
Textual Visual Exec.Rate Low-Level High-Level
93.17 73.02 77.61
97.00 78.06 82.69
96.83 77.80 81.90
96.33 78.56 82.44
Table 4: Ablation studies on various reward functions.
Reward Ablation To systematically evaluate the effectiveness of our proposed multi-granularity reward systems, we conduct ablation studies on various reward functions. To isolate the impact of the reward functions from our two-stage RL strategy, all experiments in Table 4 are trained using only a single stage. The results show that merely textual reward and combining textual and visual rewards both improve the performance in both low-level and high-level scores to a new SOTA. Also, the result demonstrates that directly utilizing visual rewards on the SFT model is not as effective as the textual reward. The detailed training process is illustrated in Figure 4. However, both of them show significant performance improvement compared with the SFT model, which leads us to explore the two-stage RL training strategy.
Stage Ablation. We also conduct an ablation study on our proposed two-stage RL training strategy. In this study, we compare three different reward settings: (1) a combination of textual and visual rewards in a single stage; (2) textual rewards in the first stage and visual rewards in the second; and (3) textual rewards in the first stage followed by both textual and visual rewards in the second. As shown in Table 5, using textual rewards in the first stage followed by combined textual and visual rewards in the second yields the best performance.
Figure 5: Showcasing charts generated by MSRL compared to proprietary and open-source MLLMs. The charts produced by MSRL align well with their ground-truth ones.
RL Stage ChartMimic
Stage 1 Stage 2 Exec.Rate Low-Level High-Level
Textual + Visual 96.33 78.56 82.44
Textual Visual 96.17 77.29 82.61
Textual Textual + Visual 96.50 78.62 83.83
Table 5: The ablation study of the reward functions on various RL stages.
4.5 Visualization
To provide a qualitative perspective on performance, we showcase several illustrative examples. As visually demonstrated in Figure 5, the charts generated by MSRL exhibit a markedly higher fidelity to the ground truth when compared with leading open-source MLLMs, even GPT-4o. In particular, the data values and layouts of charts generated by MSRL are significantly more precise.
5 Conclusion
In this work, we investigate the performance plateau of Supervised Fine-Tuning (SFT) in chart-to-code generation and propose a Multimodal Structured Reinforcement Learning (MSRL) strategy to overcome it. We begin by constructing a 3M-pair training corpus from real-world tables to systematically demonstrate that merely increasing SFT data leads to diminishing returns. To overcome this limitation, we introduce the MSRL strategy. It optimizes the saturated SFT model through a two-stage training process, guided by a novel multi-granularity reward function that assesses both textual and visual correctness. The textual reward employs a rule-based method to verify code accuracy across five key aspects, while the visual reward utilizes a render-and-compare mechanism, where an MLLM scores the similarity between the generated chart and the ground truth. Our experiments demonstrate that our model achieves new state-of-the-art performance on standard benchmarks, surpassing all open-source models and rivaling powerful proprietary ones. We make significant contributions to the field, including the curation of a large-scale dataset for a definitive empirical analysis of the SFT performance bottleneck, and the development of the MSRL framework.